# COBRA: A Continual Learning Approach to Vision-Brain Understanding

### v1

@@ 4.1 Dataset and Protocol
Protocol. We propose two different CL scenarios. First, in the (3,4)-(6,8)-(1,2)-(5,7) setup, training begins with subjects (3,4), followed by subjects (6,8), (1,2), and (5,7) sequentially, for a total of four training steps. Second, in the (3,4,6,8)-(1,2,5,7) setup, the model is trained on subjects (3, 4, 6, 8) first, then (1, 2, 5, 7), for a total of two training steps. To evaluate the CL performance, we take the final model in each training scenario and test it on the test set of subjects (3,4,6,8) and (1,2,5,7). We also evaluate the performance of these models on all subjects.
Metrics. Similar to prior studies [10, 80, 31, 69, 3, 24], we adopt the Structural Similarity Index (SSIM) and Contrastive Language-Image Pre-Training (CLIP) score to evaluate the low-level and high-level reconstruction results.
@@ 4.2 Rehearsal-free Continual Learning
In this benchmark, we do not use any samples from the previous steps to train at the current step. The performance is presented in the Table 2. Without continual learning (W/o CL), performance drops significantly to 0.207 for the Structural Similarity Index (SSIM) and 59.8% for CLIP. Our COBRA approach achieves the scores of 0.320 for SSIM and 93.2% for CLIP, on average. It outperforms previous methods such as PLOP [12] and LwF [30] by 0.04 and 0.073 for SSIM, and 17.2% and 25.8% for CLIP, respectively. In the (3,4,6,8)-(1,2,5,7) setup, we observed similar results. COBRA achieves the scores of 0.329 for SSIM and 92.9 for CLIP, on average. It outperforms PLOP by 0.037 for SSIM and 14.8% for CLIP and LwF by 0.08 for SSIM and 28.6% for CLIP. Interestingly, we found that our COBRA can cooperate with the PLOP [12] method. By doing so, COBRA achieves higher performance, i.e., 0.01 for SSIM and 1% for CLIP in the 4-step scenario, and 0.004 for SSIM and 0.8% for CLIP in the 2-step scenario.
| Method | (3,4)-(6,8)-(1,2)-(5,7) (4 steps) | (3,4,6,8)-(1,2,5,7) (2 steps) | ||||||||||
| (3,4,6,8) | (1,2,5,7) | all | (3,4,6,8) | (1,2,5,7) | all | |||||||
| SSIM | CLIP | SSIM | CLIP | SSIM | CLIP | SSIM | CLIP | SSIM | CLIP | SSIM | CLIP | |
| W/o CL | 0.197 | 58.9 | 0.216 | 60.6 | 0.207 | 59.8 | 0.197 | 59.7 | 0.238 | 61.1 | 0.218 | 60.4 |
| LwF [30] | 0.233 | 69.1 | 0.277 | 65.6 | 0.255 | 67.4 | 0.235 | 63.8 | 0.263 | 64.7 | 0.249 | 64.3 |
| PLOP [12] | 0.268 | 75.8 | 0.308 | 76.2 | 0.288 | 76.0 | 0.272 | 76.8 | 0.311 | 79.1 | 0.292 | 78.0 |
| COBRA (Ours) | 0.311 | 93.2 | 0.344 | 93.1 | 0.328 | 93.2 | 0.316 | 92.3 | 0.342 | 93.4 | 0.329 | 92.9 |
| COBRA (Ours)∗ | 0.325 | 94.4 | 0.351 | 94.8 | 0.338 | 94.2 | 0.322 | 93.5 | 0.345 | 93.8 | 0.333 | 93.7 |
| COBRA (Ours)† | 0.336 | 96.9 | 0.365 | 97.2 | 0.351 | 97.1 | 0.336 | 96.9 | 0.365 | 97.2 | 0.351 | 97.1 |

### v3

@@ 4.1 Dataset and Protocol
Protocol. We propose two different CL scenarios. First, in the (3,4)-(6,8)-(1,2)-(5,7) setup, training begins with subjects (3,4), followed by subjects (6,8), (1,2), and (5,7) sequentially, for a total of four training steps. Second, in the (3, 4, 6, 8)-(1, 2, 5, 7) setup, the model is trained on subjects (3, 4, 6, 8) first, followed by subjects (1, 2, 5, 7), for a total of two training steps. To evaluate the CL performance, we use the final model from each training scenario and test it on the test sets of subjects (3, 4, 6, 8) and (1, 2, 5, 7). We begin with the (3,4,6,8) groups (incomplete subjects), then add (1,2,5,7) (complete subjects) to simulate a realistic scenario where earlier data is limited, and richer data becomes available over time. In addition, we report performance separately on the (3,4,6,8) and (1,2,5,7) groups to enable fair comparisons between incomplete and complete data cohorts. This also allows for direct performance comparison with prior works, which often report results on the (1,2,5,7) set. We also evaluate the performance of these models on all subjects.
Metrics. For vision-brain reconstruction task, similar to prior studies [67, 68, 69, 70, 71, 72], we adopt the Structural Similarity Index (SSIM) and Contrastive Language-Image Pre-Training (CLIP) score to evaluate the low-level and high-level reconstruction results. For brain caption task, similar to prior studies [26], we employ BLEU-1, BLEU-2, BLEU-3, BLEU-4, METEOR, ROUGE, CIDEr, SPICE, CLIP-S, and RefCLIP-S to evaluate the caption results.
@@ 4.2 Rehearsal-free Continual Learning
4.2.1 Rehearsal-free Vision-Brain Reconstruction Task
@@
In this benchmark, we do not use any samples from the previous steps to train at the current step. The performance is presented in the Table 2. Without continual learning (W/o CL), performance drops significantly to 0.207 for the Structural Similarity Index (SSIM) and 59.8% for CLIP. Our COBRA approach achieves the scores of 0.320 for SSIM and 93.2% for CLIP, on average. It outperforms previous methods, such as PLOP [81] and LwF [82], by 0.04 and 0.073 for SSIM, and by 17.2% and 25.8% for CLIP, respectively. In the (3,4,6,8)-(1,2,5,7) setup, we observed similar results. COBRA achieves the scores of 0.329 for SSIM and 92.9 for CLIP, on average. It outperforms PLOP by 0.037 for SSIM and 14.8% for CLIP and LwF by 0.08 for SSIM and 28.6% for CLIP. Interestingly, we found that our COBRA can cooperate with the PLOP [81] method. By doing so, COBRA achieves higher performance, i.e., 0.01 for SSIM and 1% for CLIP in the 4-step scenario, and 0.004 for SSIM and 0.8% for CLIP in the 2-step scenario.
| (3,4)-(6,8)-(1,2)-(5,7) (4 steps) | ||||||
| Method | (3,4,6,8) | (1,2,5,7) | all | |||
| SSIM | CLIP | SSIM | CLIP | SSIM | CLIP | |
| W/o CL | 0.1970.005 | 58.90.6 | 0.2160.006 | 60.60.7 | 0.2070.005 | 59.80.6 |
| LwF [82] | 0.2330.006 | 69.10.5 | 0.2770.007 | 65.60.5 | 0.2550.007 | 67.40.4 |
| PLOP [81] | 0.2680.005 | 75.80.4 | 0.3080.005 | 76.20.4 | 0.2880.005 | 76.00.4 |
| COBRA | 0.3110.004 | 93.20.3 | 0.3440.004 | 93.10.3 | 0.3280.004 | 93.20.3 |
| COBRA∗ | 0.3250.003 | 94.40.2 | 0.3510.003 | 94.80.2 | 0.3380.003 | 94.20.2 |
| COBRA† | 0.3360.002 | 96.90.2 | 0.3650.002 | 97.20.2 | 0.3510.002 | 97.10.2 |
| (3,4,6,8)-(1,2,5,7) (2 steps) | ||||||
| Method | (3,4,6,8) | (1,2,5,7) | all | |||
| SSIM | CLIP | SSIM | CLIP | SSIM | CLIP | |
| W/o CL | 0.1970.005 | 59.70.6 | 0.2380.007 | 61.10.6 | 0.2180.006 | 60.40.6 |
| LwF [82] | 0.2350.006 | 63.80.5 | 0.2630.006 | 64.70.4 | 0.2490.006 | 64.30.4 |
| PLOP [81] | 0.2720.005 | 76.80.4 | 0.3110.006 | 79.10.5 | 0.2920.005 | 78.00.5 |
| COBRA | 0.3160.004 | 92.30.3 | 0.3420.004 | 93.40.3 | 0.3290.004 | 92.90.3 |
| COBRA∗ | 0.3220.003 | 93.50.2 | 0.3450.003 | 93.80.2 | 0.3330.003 | 93.70.2 |
| COBRA† | 0.3360.002 | 96.90.2 | 0.3650.002 | 97.20.2 | 0.3510.002 | 97.10.2 |
4.2.2 Rehearsal-free Brain Captioning Task
Building on the models trained for the Vision-Brain Reconstruction task described in Section 4.2.1 and following the approach in [26], we employ a linear projector or adapter to map the fMRI feature representations into the language embedding space. These transformed features are then passed to a large language model (LLM), specifically Shikra [83], to generate captions corresponding to brain signals. We evaluate the performance of rehearsal-free continual learning for this brain captioning task across multiple metrics: BLEU-1 and BLEU-2 in Table 3, BLEU-3 and BLEU-4 in Table 4, METEOR and ROUGE in Table 5, CIDEr and SPICE in Table 6, and CLIP-S and RefCLIP-S in Table 7. Overall, the performance trends are consistent with those observed in the vision-brain reconstruction task. Across most metrics and both continual learning settings (4-step and 2-step), our proposed method, COBRA, consistently outperforms competing approaches, including LwF [82], PLOP [81], and the baseline without continual learning by a substantial margin. These results further highlight the robustness and effectiveness of COBRA across diverse vision-brain understanding tasks.
| (3,4)-(6,8)-(1,2)-(5,7) (4 steps) | ||||||
| Method | (3,4,6,8) | (1,2,5,7) | all | |||
| BLEU-1 | BLEU-2 | BLEU-1 | BLEU-2 | BLEU-1 | BLEU-2 | |
| W/o CL | 47.20.4 | 31.20.3 | 49.10.4 | 32.90.3 | 48.10.4 | 32.00.3 |
| LwF [82] | 51.00.3 | 34.00.3 | 52.30.3 | 35.10.3 | 51.60.3 | 34.60.3 |
| PLOP [81] | 53.30.3 | 35.70.3 | 54.60.3 | 36.90.3 | 53.90.3 | 36.30.3 |
| COBRA | 58.60.2 | 39.80.2 | 59.40.2 | 40.50.2 | 59.00.2 | 40.20.2 |
| COBRA∗ | 59.20.2 | 40.40.2 | 60.20.2 | 41.20.2 | 59.70.2 | 40.90.2 |
| COBRA† | 62.60.2 | 41.70.2 | 62.00.2 | 43.10.2 | 62.30.2 | 42.40.2 |
| (3,4,6,8)-(1,2,5,7) (2 steps) | ||||||
| Method | (3,4,6,8) | (1,2,5,7) | all | |||
| BLEU-1 | BLEU-2 | BLEU-1 | BLEU-2 | BLEU-1 | BLEU-2 | |
| W/o CL | 47.50.4 | 31.60.3 | 48.90.4 | 32.70.3 | 48.20.4 | 32.20.3 |
| LwF [82] | 51.20.3 | 34.20.3 | 52.60.3 | 35.50.3 | 51.90.3 | 34.90.3 |
| PLOP [81] | 53.60.3 | 36.00.3 | 54.80.3 | 37.10.3 | 54.20.3 | 36.60.3 |
| COBRA | 59.80.2 | 41.00.2 | 60.80.2 | 41.80.2 | 60.30.2 | 41.40.2 |
| COBRA∗ | 60.60.2 | 41.80.2 | 61.50.2 | 42.60.2 | 61.10.2 | 42.20.2 |
| COBRA† | 62.60.2 | 41.70.2 | 62.00.2 | 43.10.2 | 62.30.2 | 42.40.2 |
| (3,4)-(6,8)-(1,2)-(5,7) (4 steps) | ||||||
| Method | (3,4,6,8) | (1,2,5,7) | all | |||
| BLEU-3 | BLEU-4 | BLEU-3 | BLEU-4 | BLEU-3 | BLEU-4 | |
| W/o CL | 19.30.4 | 11.60.3 | 20.10.5 | 12.10.2 | 19.70.3 | 11.90.4 |
| LwF [82] | 21.70.3 | 13.30.2 | 22.60.4 | 13.90.3 | 22.10.2 | 13.60.4 |
| PLOP [81] | 23.00.2 | 14.30.5 | 24.10.4 | 15.10.3 | 23.50.3 | 14.70.3 |
| COBRA | 26.10.4 | 16.40.3 | 27.20.5 | 17.10.3 | 26.70.4 | 16.80.2 |
| COBRA∗ | 26.70.5 | 16.90.4 | 27.80.3 | 17.50.3 | 27.30.4 | 17.20.3 |
| COBRA† | 29.60.3 | 19.10.2 | 31.00.3 | 20.50.2 | 30.30.3 | 19.80.2 |
| (3,4,6,8)-(1,2,5,7) (2 steps) | ||||||
| Method | (3,4,6,8) | (1,2,5,7) | all | |||
| BLEU-3 | BLEU-4 | BLEU-3 | BLEU-4 | BLEU-3 | BLEU-4 | |
| W/o CL | 19.50.3 | 11.80.4 | 20.40.2 | 12.40.5 | 20.00.4 | 12.10.2 |
| LwF [82] | 22.00.3 | 13.50.2 | 23.00.4 | 14.10.3 | 22.60.5 | 13.90.3 |
| PLOP [81] | 23.30.4 | 14.60.2 | 24.50.3 | 15.40.4 | 23.90.2 | 15.00.5 |
| COBRA | 27.40.3 | 17.30.3 | 28.30.2 | 18.10.2 | 27.90.4 | 17.70.4 |
| COBRA∗ | 28.00.5 | 17.90.4 | 28.90.3 | 18.60.3 | 28.40.4 | 18.30.3 |
| COBRA† | 29.60.3 | 19.10.2 | 31.00.3 | 20.50.2 | 30.30.3 | 19.80.2 |
| (3,4)-(6,8)-(1,2)-(5,7) (4 steps) | ||||||
| Method | (3,4,6,8) | (1,2,5,7) | all | |||
| METEOR | ROUGE | METEOR | ROUGE | METEOR | ROUGE | |
| W/o CL | 14.20.4 | 36.50.2 | 15.10.5 | 37.30.3 | 14.60.3 | 36.90.4 |
| LwF [82] | 15.80.2 | 38.10.3 | 16.60.4 | 39.00.5 | 16.20.3 | 38.60.2 |
| PLOP [81] | 17.10.5 | 39.40.4 | 17.90.2 | 40.30.2 | 17.50.4 | 39.90.3 |
| COBRA | 18.80.4 | 42.00.3 | 19.70.3 | 42.80.4 | 19.20.4 | 42.40.3 |
| COBRA∗ | 19.30.2 | 42.50.5 | 20.30.3 | 43.30.3 | 19.80.4 | 42.90.4 |
| COBRA† | 21.70.3 | 44.10.2 | 22.90.3 | 45.40.2 | 22.30.3 | 44.80.2 |
| (3,4,6,8)-(1,2,5,7) (2 steps) | ||||||
| Method | (3,4,6,8) | (1,2,5,7) | all | |||
| METEOR | ROUGE | METEOR | ROUGE | METEOR | ROUGE | |
| W/o CL | 14.50.3 | 36.80.3 | 15.40.5 | 37.60.4 | 15.00.2 | 37.20.2 |
| LwF [82] | 16.30.4 | 38.40.2 | 17.20.5 | 39.30.3 | 16.80.3 | 38.90.5 |
| PLOP [81] | 17.60.2 | 39.70.4 | 18.50.4 | 40.60.3 | 18.10.2 | 40.20.5 |
| COBRA | 20.00.3 | 43.10.5 | 20.90.2 | 43.90.2 | 20.50.4 | 43.50.3 |
| COBRA∗ | 20.50.4 | 43.60.3 | 21.40.4 | 44.40.5 | 21.00.3 | 44.00.4 |
| COBRA† | 21.70.3 | 44.10.2 | 22.90.3 | 45.40.2 | 22.30.3 | 44.80.2 |
| (3,4)-(6,8)-(1,2)-(5,7) (4 steps) | ||||||
| Method | (3,4,6,8) | (1,2,5,7) | all | |||
| CIDEr | SPICE | CIDEr | SPICE | CIDEr | SPICE | |
| W/o CL | 35.20.4 | 6.80.3 | 37.50.5 | 7.30.4 | 36.40.2 | 7.10.4 |
| LwF [82] | 40.80.3 | 8.10.5 | 43.00.2 | 8.60.3 | 41.90.4 | 8.40.2 |
| PLOP [81] | 47.60.3 | 8.80.4 | 50.10.4 | 9.50.2 | 48.90.3 | 9.20.5 |
| COBRA | 56.40.5 | 11.20.4 | 58.70.3 | 11.90.4 | 57.60.2 | 11.60.3 |
| COBRA∗ | 57.30.3 | 11.60.2 | 59.80.4 | 12.30.5 | 58.60.3 | 12.00.4 |
| COBRA† | 62.00.4 | 13.20.3 | 64.40.4 | 13.90.3 | 63.20.4 | 13.50.3 |
| (3,4,6,8)-(1,2,5,7) (2 steps) | ||||||
| Method | (3,4,6,8) | (1,2,5,7) | all | |||
| CIDEr | SPICE | CIDEr | SPICE | CIDEr | SPICE | |
| W/o CL | 36.00.5 | 7.00.2 | 38.20.4 | 7.60.3 | 37.10.3 | 7.30.5 |
| LwF [82] | 42.20.4 | 8.40.3 | 44.60.2 | 9.00.2 | 43.40.5 | 8.70.4 |
| PLOP [81] | 49.30.2 | 9.20.4 | 52.00.5 | 9.90.3 | 50.70.4 | 9.60.3 |
| COBRA | 59.10.4 | 12.10.2 | 61.30.3 | 12.80.5 | 60.20.3 | 12.40.4 |
| COBRA∗ | 60.30.5 | 12.40.3 | 62.70.3 | 13.20.4 | 61.50.2 | 12.80.3 |
| COBRA† | 62.00.4 | 13.20.3 | 64.40.4 | 13.90.3 | 63.20.4 | 13.50.3 |
| (3,4)-(6,8)-(1,2)-(5,7) (4 steps) | ||||||
| Method | (3,4,6,8) | (1,2,5,7) | all | |||
| CLIPS | RefCLIPS | CLIPS | RefCLIPS | CLIPS | RefCLIPS | |
| W/o CL | 56.80.3 | 62.30.4 | 58.10.5 | 63.60.2 | 57.50.4 | 62.90.3 |
| LwF [82] | 59.30.2 | 64.10.3 | 60.50.4 | 65.30.3 | 59.90.5 | 64.70.2 |
| PLOP [81] | 61.00.4 | 65.80.2 | 62.40.3 | 67.00.5 | 61.70.3 | 66.40.4 |
| COBRA | 65.30.5 | 70.10.4 | 66.60.3 | 71.20.2 | 66.00.3 | 70.60.5 |
| COBRA∗ | 66.00.2 | 70.80.5 | 67.40.4 | 72.00.3 | 66.70.3 | 71.40.2 |
| COBRA† | 69.30.3 | 73.60.2 | 70.10.3 | 75.00.2 | 70.20.3 | 74.30.2 |
| (3,4,6,8)-(1,2,5,7) (2 steps) | ||||||
| Method | (3,4,6,8) | (1,2,5,7) | all | |||
| CLIPS | RefCLIPS | CLIPS | RefCLIPS | CLIPS | RefCLIPS | |
| W/o CL | 57.70.5 | 63.10.3 | 59.00.4 | 64.20.2 | 58.30.3 | 63.60.5 |
| LwF [82] | 60.20.3 | 64.90.3 | 61.50.5 | 66.10.4 | 60.80.2 | 65.50.2 |
| PLOP [81] | 62.00.3 | 66.50.4 | 63.50.2 | 67.70.5 | 62.80.4 | 67.10.3 |
| COBRA | 67.00.2 | 71.50.2 | 68.20.4 | 72.70.5 | 67.60.3 | 72.10.3 |
| COBRA∗ | 68.00.3 | 72.50.4 | 69.20.3 | 73.60.3 | 68.60.4 | 73.10.2 |
| COBRA† | 69.30.3 | 73.60.2 | 70.10.3 | 75.00.2 | 70.20.3 | 74.30.2 |
@@ 4.3 Rehearsal-based Continual Learning
4.3.1 Rehearsal-based Vision-Brain Reconstruction Task